# Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation

### v2

@@ Abstract
Model merging aims to integrate multiple task-specific models into a unified model that inherits the capabilities of the task-specific models, without additional training. Existing model merging methods often lack consideration of the varying contribution ratios of different task-specific models to the final merged model. In this paper, we propose Mixup Model Merge (M3), a simple yet effective method inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique. M3 performs randomized linear interpolation in parameter space between two task-specific LLMs, where interpolation coefficients are sampled from a Beta distribution to explore diverse contribution ratios. This controllable randomness allows M3 to outperform standard equal-ratio merging by discovering better contribution ratio combinations. Extensive experiments show that M3 significantly (1) improves merged LLM performance across tasks, (2) enhances out-of-distribution and adversarial robustness, and (3) outperforms the positive effects of the sparsification method DARE on model merging and can be further combined with DARE to achieve superior results. By tuning the Beta distribution’s shape parameters, (4) M3 balances exploration efficiency and diversity in contribution ratios. The code is available at https://github.com/MLGroupJLU/MixupModelMerge.
Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation
Yue Zhou1  Yi Chang1,2,3  Yuan Wu1111Corresponding authors 1School of Artificial Intelligence, Jilin University 2Engineering Research Center of Knowledge-Driven Human-Machine Intelligence, MOE, China 3International Center of Future Science, Jilin University yuezhou24@mails.jlu.edu.cn, yichang@jlu.edu.cn, yuanwu@jlu.edu.cn
@@ 1 Introduction
In the field of Natural Language Processing (NLP), the emergence of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; OpenAI, 2023; Chowdhery et al., 2023) represents a revolutionary breakthrough. With their remarkable capabilities, these models have demonstrated outstanding performance across various tasks (Jiao et al., 2023; Chang et al., 2024b; Nam et al., 2024; Xing, 2024; Guo et al., 2024), significantly advancing NLP technologies.
(a) Performance of the merged models obtained through Task Arithmetic and TIES-Merging with/without M3.
(b) (Left) Performance of merged models with/without M3 on OOD datasets. (Right) Adversarial robustness (PDR) of merged models with/without M3.
Figure 1: (a) M3 significantly boosts the model merging performance. (b) OOD robustness, and adversarial robustness of the merged models.
Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al., 2021; Ding et al., 2023; Xia et al., 2024). However, SFT requires substantial computational resources and long training times (Brown et al., 2020; Chang et al., 2024a). To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al., 2024; Akiba et al., 2024). However, Existing model merging methods often adopt identical contribution ratios for different task-specific models in the merging process (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024), or rely on a few manually predefined contribution ratio values (Lin et al., 2023). This restricts the exploration of how varying contribution ratios influence the final merged model, and limits the ability to fully explore the resulting parameter space. Consequently, the potential of the merged model is significantly constrained.
To address the above issue, we propose a novel approach, Mixup Model Merge (M3), for exploring appropriate contribution ratios among task-specific models. This method is inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique (Zhang, 2017). In contrast to conventional Mixup, M3 performs linear interpolation in the model parameter space, where the interpolation coefficients are randomly sampled from a Beta distribution. Specifically, this method controls the contribution ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$. By adjusting the parameter $\alpha$, we can precisely control the distribution of $\lambda_{m}$, enabling M3 to balance the diversity of $\lambda_{m}$ values with the efficiency of discovering suitable contribution ratios. In the parameter space, our method explores the family of models formed along the linear path between two task-specific LLMs by randomly sampling interpolation coefficients. This allows us to identify suitable interpolation coefficients (i.e., contribution ratios), thereby obtaining a merged model with superior performance. As shown in Figure 2, the implementation of M3 is similar to the process of proportionally mixing magical potions in Harry Potter.
We conducted extensive experiments with three homologous task-specific fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively. Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang, 2017), we further performed comprehensive evaluations on LiveBench (White et al., 2024) and PromptBench (Zhu et al., 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness (Wang et al., 2023; Zhu et al., 2023) of merged models. The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), and enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)).
@@ 2 Related Works
Model Merging
Model merging is a technique that integrates the parameters of multiple models to create a unified model with enhanced or diverse capabilities (Wortsman et al., 2022; Ilharco et al., 2022; Matena and Raffel, 2022; Jin et al., 2022; Yadav et al., 2023; Yu et al., 2024; Lin et al., 2024). Task arithmetic (Ilharco et al., 2022) leverages task vectors for model merging through arithmetic operations, incorporating a predefined scaling term to weight the contribution of different models. Fisher Merging (Matena and Raffel, 2022) performs parameter fusion by applying weights derived from the Fisher information matrix (Fisher, 1922), resulting in more precise parameter integration. TIES-Merging (Yadav et al., 2023) addresses task conflicts by removing low-magnitude parameters, resolving sign disagreements, and merging only the parameters that align with the final agreed-upon sign. In (Yu et al., 2024), it is found that LLMs can enhance their capabilities through model merging. Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al., 2022), significantly improving the performance of various model merging techniques. DELLA (Deep et al., 2024) is a novel model merging technique that integrates a new pruning strategy called MAGPRUNE, which samples delta parameters based on their magnitudes. MAGPRUNE demonstrates improvements over existing methods such as DARE and TIES.
@@ 3.1 Mixup Model Merge
Inspired by the random interpolation strategy of Mixup (Zhang, 2017), we extend the idea from input space to parameter space, proposing Mixup Model Merge (M3).
@@
$\theta_{M}=\lambda_{m}\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\theta_{\text{SFT}}^{t_{2}},\quad\lambda_{m}\in(0,1)$
@@
$\delta_{t_{i}}=\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}},\quad i\in\{1,2\}$
@@
$\theta_{M}=\theta_{\text{PRE}}+\lambda_{m}\delta_{t_{1}}+(1-\lambda_{m})\delta_{t_{2}}$
@@
$\mathcal{F}_{M}=\{\theta_{M}^{\lambda_{m}}\mid\lambda_{m}\in(0,1)\}$
@@
Linear Interpolation Coefficient
@@
Comparison with Grid Search
As shown in Table 1, we compare the performance of M3 with that of grid search. Grid search requires manually specifying multiple fixed and discrete interpolation coefficients (e.g., $\lambda=0.1,0.2,\ldots,0.9$). While this approach offers fine-grained control, it lacks the ability to explore continuous values. Achieving finer resolution (e.g., identifying values beyond the first decimal place) requires a smaller step size, which significantly reduces search efficiency.
In contrast, our method enables more efficient and diverse exploration of continuous interpolation coefficients. For example, our approach can sample values such as $\lambda_{m}=0.62156685$, which are difficult to obtain with coarse grid search. Empirically, we find that the model’s performance is often sensitive to the second decimal place of $\lambda$. To achieve such granularity, grid search would require a step size of 0.01, resulting in a prohibitively large number of evaluations.
Our method, by comparison, typically requires only a single pass over a small set of hyperparameter values for $\alpha$ (e.g., $\alpha\in\{0.2,0.4,0.5,1,2,3,5\}$). In cases where a suitable $\lambda_{m}$ is not found in the first pass, we perform a few additional samplings under a selected $\alpha$ value—no more than three times in our experiments. In Table 2, we report the specific $\lambda_{m}$ values that lead to performance improvements, along with the corresponding distribution shape parameter $\alpha$ and the total number of sampling iterations.
Method Distribution of $\lambda_{m}$ Continuity of $\lambda_{m}$ Efficiency Diversity Controllability
Grid Search Fixed discrete values (e.g., 0.1–0.9) No Low Medium High
M3+Beta(1,1) Uniform over (0,1) Yes Medium High Medium
M3+Beta(5,5) Concentrated near the center (0.5) Yes High Low High
M3+Beta(0.2,0.2) Concentrated near the endpoints (0 or 1) Yes High Low High
Table 1: Comparison of interpolation coefficient $\lambda_{m}$ sampling strategies.
M3 Implementation
M3 is a plug-and-play approach that can be integrated into existing representative model merging methods to enhance the performance of the resulting merged models. As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al., 2022) and Task Arithmetic (Ilharco et al., 2022).
@@
$\theta_{M}=\frac{1}{2}\left(\theta_{\text{SFT}}^{t_{1}}+\theta_{\text{SFT}}^{t_{2}}\right)$ (1)
@@
$\theta_{M}$ $=\theta_{\text{PRE}}+\lambda\cdot(\delta^{t_{1}}+\delta^{t_{2}})$
$=\theta_{\text{PRE}}+\lambda\cdot\sum_{i=1}^{2}(\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}}),$ (2)
@@
$\lambda_{m}\in[0,1],\quad\lambda_{m}\sim\mathrm{Beta}(\alpha,\alpha),$
@@ 3.2 Theoretical Analysis
According to the theoretical and empirical findings of Neyshabur et al. (2020), fine-tuning models from the same pretrained initialization can guide them into a shared, flat, and connected low-loss basin. As a result, the linear interpolation path between such fine-tuned models remains within this low-loss basin and does not encounter significant performance barriers. In some cases, models along the interpolation path even outperform the endpoints.
@@
$f_{i}(\lambda_{m})=\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i}.$ (6)
@@ 4.1 Experimental Setup
Task-Specific Fine-Tuned LLMs and Datasets
Following the experimental setup given in Yu et al. (2024), we select three task-specific fine-tuned LLMs: WizardLM-13B (Xu et al., 2024), WizardMath-13B (Luo et al., 2023), and llama-2-13b-code-alpaca (Chaudhary, 2023), all of which use Llama-2-13b (Touvron et al., 2023) as the pre-trained backbone. These models are respectively designed for instruction-following, mathematical reasoning, and code generation tasks. To evaluate the instruction-following task we use AlpacaEval (Li et al., 2023). For testing mathematical reasoning task, we employ GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For estimating the performance of code-generating task, we use HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021). More details of these LLMs and datasets can be found in Appendix A.1.
Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness
To evaluate OOD robustness, we evaluate merged models including math & code, LM & math, and LM & code on instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al., 2024), respectively. More details on OOD benchmarks are given in Appendix A.3.
We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al., 2024) to evaluate the robustness of LLMs against adversarial prompts. Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al., 2018), BERTAttack (word-level) (Li et al., 2020), and StressTest (sentence-level) (Naik et al., 2018). The evaluation is conducted on two datasets supported by PromptBench: SST2 (sentiment analysis) (Socher et al., 2013) and CoLA (grammatical correctness) (Warstadt, 2019). For more details on PromptBench and attack methods, please refer to Appendix A.4.
Evaluation Metrics
@@
Unless otherwise specified, the remaining details of the model merging experiments follow Yu et al. (2024). For a detailed description of the hyperparameter settings used in the representative existing model merging methods, please refer to Appendix A.2. All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
4.2 Merging Task-Specific Fine-Tuned LLMs
Merging Methods Models Use M3 (Ours) Sampling Details Instruction Following Mathematical Reasoning Code Generating Avg.
$\lambda_{m}$ $\alpha$ AlpacaEval GSM8K MATH HumanEval MBPP
- LM - - - 45.14 2.20 0.04 36.59 34.00 -
Math - - - - 64.22 14.02 - - -
Code - - - - - - 23.78 27.60 -
Average Merging LM & Math No - - 45.28 66.34 13.40 - - 41.67
Yes 0.35 2 44.40 66.26 13.80 - - 41.48
LM & Code No - - 36.60 - - 29.88 32.00 32.82
Yes 0.84 2 43.91$\uparrow$ - - 37.20$\uparrow$ 34.40 38.50$\uparrow$
Math & Code No - - - 56.17 10.28 8.53 8.20 20.80
Yes 0.92 0.4 - 63.61$\uparrow$ 14.02$\uparrow$ 8.54 19.20$\uparrow$ 26.34$\uparrow$
Task Arithmetic LM & Math No - - 45.78 66.34 13.40 - - 41.84
Yes 0.34 2 41.65 66.34 13.74 - - 40.58
LM & Code No - - 44.64 - - 32.93 33.60 37.06
Yes 0.99 0.4 46.64 - - 35.37 33.80 38.60
Math & Code No - - - 64.67 13.98 8.54 8.60 23.95
Yes 0.98 0.5 - 63.53 13.94 7.93 19.00$\uparrow$ 26.1
TIES- Merging LM & Math No - - 38.63 14.56 2.12 - - 18.44
Yes 0.62 1 38.73 18.57$\uparrow$ 2.48 - - 19.92
LM & Code No - - 41.85 - - 0.0 0.0 13.95
Yes 0.84 0.5 44.96 - - 25.61$\uparrow$ 30.80$\uparrow$ 33.79$\uparrow$
Math & Code No - - - 64.67 13.68 9.15 22.60 27.57
Yes 0.63 3 - 64.75 14.16 9.76 21.4 27.52
DARE LM & Math No - - 49.00 66.64 13.2 - - 42.95
Yes 0.38 0.4 44.90 67.32 13.74 - - 41.98
LM & Code No - - 41.47 - - 35.98 33.00 36.82
Yes 0.99 0.5 45.20$\uparrow$ - - 35.98 35.20 38.79
Math & Code No - - - 65.05 10.37 9.15 9.80 23.59
Yes 0.98 0.5 - 65.13 14.32$\uparrow$ 8.54 18.00$\uparrow$ 26.50$\uparrow$
Performance Gains of M3 Across Merging Methods and Tasks
In addition, we also compare our proposed method M3 with existing model merging methods. The details are in Appendix G.
Mitigating Suboptimal Fine-Tuning with M3
4.3 Model Robustness
Out-of-distribution robustness
To ensure that the evaluation datasets are as representative as possible of OOD data, we select datasets with sufficiently recent release dates and ensure they cover domains that fine-tuned LLMs have not been specifically trained on. Consequently, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing. As illustrated in Figure 4, M3 consistently enhances the performance of merged models—Math & Code, LM & Math, and LM & Code—on OOD datasets. Specifically, when Task Arithmetic is combined with M3, the Math & Code model demonstrates a 1.9% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 1.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model shows a significant 6% boost in accuracy on LiveBench-TypoFixing. Similarly, when Average Merging is combined with M3, the Math & Code model attains a 1.5% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 0.7% increase in pass@1 on LiveBench-Coding, and the LM & Code model exhibits a 4% enhancement in accuracy on LiveBench-TypoFixing. Finally, when TIES-Merging is applied alongside M3, the Math & Code model achieves a 1.1% improvement in win rate on LiveBench-Instruction, the LM & Math model records a 0.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model demonstrates a remarkable 14% improvement in accuracy on LiveBench-TypoFixing.
Adversarial robustness
Model Dataset Use Mixup Use Attack Metric (%) PDR (%)
Math & Code SST2 No No 57.68 38.97

### v3

@@ Abstract
Model merging aims to integrate multiple task-specific models into a unified model that inherits the capabilities of the task-specific models, without additional training. Existing model merging methods often lack consideration of the varying contribution ratios of different task-specific models to the final merged model. In this paper, we propose Mixup Model Merge (M3), a simple yet effective method inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique. M3 performs randomized linear interpolation in parameter space between two task-specific LLMs, where interpolation coefficients are sampled from a Beta distribution to explore diverse contribution ratios. This controllable randomness allows M3 to outperform standard equal-ratio merging by discovering better contribution ratio combinations. Extensive experiments show that M3 significantly (1) improves merged LLM performance across tasks, (2) enhances out-of-distribution and adversarial robustness, and (3) outperforms the positive effects of the sparsification method DARE on model merging and can be further combined with DARE to achieve superior results. By tuning the Beta distribution’s shape parameters, (4) M3 balances exploration efficiency and diversity in contribution ratios. The code is provided in the supplementary materials.
Codehttps://github.com/MLGroupJLU/MixupModelMerge
@@ 1 Introduction
In the field of Natural Language Processing (NLP), the emergence of large language models (LLMs) (Brown et al. 2020; Touvron et al. 2023; OpenAI 2023; Chowdhery et al. 2023) represents a revolutionary breakthrough. With their remarkable capabilities, these models have demonstrated outstanding performance across various tasks (Jiao et al. 2023; Chang et al. 2024; Nam et al. 2024; Xing 2024; Guo et al. 2024), significantly advancing NLP technologies.
(a) Performance of the merged models obtained through Task Arithmetic and TIES-Merging with or without M3
(b) (Left) Performance on OOD datasets. (Right) Adversarial robustness (PDR) of merged models with or without M3
Figure 1: (a) M3 significantly improves the performance of model merging. (b) OOD and adversarial robustness of merged models with or without M3
@@
Supervised Fine-Tuning (SFT) is a crucial technique for adapting LLMs to specific tasks, refining their performance by training on domain-specific data (Hu et al. 2021; Ding et al. 2023; Xia et al. 2024). However, SFT requires substantial computational resources and long training times (Brown et al. 2020; Chang, Chang, and Wu 2024). To address this challenge, Model Merging has emerged as an efficient solution, fusing the parameters of multiple fine-tuned LLMs into a unified model with diverse capabilities, without the need for additional training or computational costs (Yang et al. 2024; Akiba et al. 2024). However, Existing model merging methods often adopt identical contribution ratios for different task-specific models in the merging process (Wortsman et al. 2022; Ilharco et al. 2022; Matena and Raffel 2022; Jin et al. 2022; Yadav et al. 2023; Yu et al. 2024), or rely on a few manually predefined contribution ratio values (Lin et al. 2023). This restricts the exploration of how varying contribution ratios influence the final merged model, and limits the ability to fully explore the resulting parameter space. Consequently, the potential of the merged model is significantly constrained.
To address the above issue, we propose a novel approach, Mixup Model Merge (M3), for exploring appropriate contribution ratios among task-specific models. This method is inspired by the randomized linear interpolation strategy from the Mixup data augmentation technique (Zhang 2017). In contrast to conventional Mixup, M3 performs linear interpolation in the model parameter space, where the interpolation coefficients are randomly sampled from a Beta distribution. Specifically, this method controls the contribution ratio between two fine-tuned LLMs by randomly generating a linear interpolation ratio $\lambda_{m}$, where $\lambda_{m}\in(0,1)$ and $\lambda_{m}\sim\text{Beta}(\alpha,\alpha)$. By adjusting the parameter $\alpha$, we can precisely control the distribution of $\lambda_{m}$, enabling M3 to balance the diversity of $\lambda_{m}$ values with the efficiency of discovering suitable contribution ratios. In the parameter space, our method explores the family of models formed along the linear path between two task-specific LLMs by randomly sampling interpolation coefficients. This allows us to identify suitable interpolation coefficients (i.e., contribution ratios), thereby obtaining a merged model with superior performance.
We conducted extensive experiments with three homologous task-specific fine-tuned LLMs: WizardLM-13B (Xu et al. 2024), WizardMath-13B (Luo et al. 2023), and llama-2-13b-code-alpaca (Chaudhary 2023), which specialize in instruction following, mathematical reasoning, and code generation, respectively. Inspired by Mixup’s effectiveness in enhancing the robustness of neural networks when handling corrupted labels or adversarial examples (Zhang 2017), we further performed comprehensive evaluations on LiveBench (White et al. 2024) and PromptBench (Zhu et al. 2024) to validate the potential of M3 in improving the OOD robustness and adversarial robustness (Wang et al. 2023; Zhu et al. 2023) of merged models. The experimental results demonstrate that our proposed M3 method can significantly improve merged models’ performance across various tasks (as shown in Figure 1(a)), and enhance the OOD and adversarial robustness of the merged models (as shown in Figure 1(b)).
@@ 2 Related Works
2.1 Model Merging
Model merging is a technique that integrates the parameters of multiple models to create a unified model with enhanced or diverse capabilities (Wortsman et al. 2022; Ilharco et al. 2022; Matena and Raffel 2022; Jin et al. 2022; Yadav et al. 2023; Yu et al. 2024; Lin et al. 2024). Task arithmetic (Ilharco et al. 2022) leverages task vectors for model merging through arithmetic operations, incorporating a predefined scaling term to weight the contribution of different models. Fisher Merging (Matena and Raffel 2022) performs parameter fusion by applying weights derived from the Fisher information matrix (Fisher 1922), resulting in more precise parameter integration. TIES-Merging (Yadav et al. 2023) addresses task conflicts by removing low-magnitude parameters, resolving sign disagreements, and merging only the parameters that align with the final agreed-upon sign. In (Yu et al. 2024), it is found that LLMs can enhance their capabilities through model merging. Additionally, it introduces DARE, a method for sparsifying the delta parameters of the model (Ilharco et al. 2022), significantly improving the performance of various model merging techniques. DELLA (Deep, Bhardwaj, and Poria 2024) is a novel model merging technique that integrates a new pruning strategy called MAGPRUNE, which samples delta parameters based on their magnitudes. MAGPRUNE demonstrates improvements over existing methods such as DARE and TIES.
@@ 3.1 Mixup Model Merge
Inspired by the random interpolation strategy of Mixup (Zhang 2017), we extend the idea from input space to parameter space, proposing Mixup Model Merge (M3).
@@
$\theta_{M}=\lambda_{m}\theta_{\text{SFT}}^{t_{1}}+(1-\lambda_{m})\theta_{\text{SFT}}^{t_{2}},\quad\lambda_{m}\in(0,1),$ (1)
@@
$\delta_{t_{i}}=\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}},\quad i\in\{1,2\},$ (2)
@@
$\theta_{M}=\theta_{\text{PRE}}+\lambda_{m}\delta_{t_{1}}+(1-\lambda_{m})\delta_{t_{2}},$ (3)
@@
$\mathcal{F}_{M}=\{\theta_{M}^{\lambda_{m}}\mid\lambda_{m}\in(0,1)\},$ (4)
@@
Linear Interpolation Coefficient.
@@
In Table 1, we report the specific $\lambda_{m}$ values that lead to performance improvements, along with the corresponding distribution shape parameter $\alpha$.
M3 Implementation.
M3 is a plug-and-play approach that can be integrated into existing representative model merging methods to enhance the performance of the resulting merged models. As illustrative examples, we consider two widely used merging methods: Average Merging (Wortsman et al. 2022) and Task Arithmetic (Ilharco et al. 2022).
@@
$\theta_{M}=\frac{1}{2}\left(\theta_{\text{SFT}}^{t_{1}}+\theta_{\text{SFT}}^{t_{2}}\right),$ (5)
@@
$\theta_{M}$ $=\theta_{\text{PRE}}+\lambda\cdot(\delta^{t_{1}}+\delta^{t_{2}}),$ (6)
$=\theta_{\text{PRE}}+\lambda\cdot\sum_{i=1}^{2}(\theta_{\text{SFT}}^{t_{i}}-\theta_{\text{PRE}}),$
@@
$\lambda_{m}\in[0,1],\quad\lambda_{m}\sim\mathrm{Beta}(\alpha,\alpha),$ (9)
@@
Merging
Methods
Models
Use M3
Sampling
Details
Instruction
Following
Mathematical
Reasoning
Code Generating
Avg.
$\lambda_{m}$ $\alpha$ AlpacaEval GSM8K MATH HumanEval MBPP
- LM - - - 45.14 2.20 0.04 36.59 34.00 -
Math - - - - 64.22 14.02 - - -
Code - - - - - - 23.78 27.60 -
Average
Merging
LM
& Math
No - - 45.28 66.34 13.40 - - 41.67
Yes 0.35 2 44.40 66.26 13.80 - - 41.48
LM
& Code
No - - 36.60 - - 29.88 32.00 32.82
[…]
Math
& Code
No - - - 56.17 10.28 8.53 8.20 20.80
[…]
Task
Arithmetic
LM
& Math
No - - 45.78 66.34 13.40 - - 41.84
Yes 0.34 2 41.65 66.34 13.74 - - 40.58
LM
& Code
No - - 44.64 - - 32.93 33.60 37.06
Yes 0.99 0.4 46.64 - - 35.37 33.80 38.60
Math
& Code
No - - - 64.67 13.98 8.54 8.60 23.95
[…]
TIES-
Merging
LM
& Math
No - - 38.63 14.56 2.12 - - 18.44
[…]
LM
& Code
No - - 41.85 - - 0.0 0.0 13.95
[…]
Math
& Code
No - - - 64.67 13.68 9.15 22.60 27.57
Yes 0.63 3 - 64.75 14.16 9.76 21.4 27.52
DARE
LM
& Math
No - - 49.00 66.64 13.2 - - 42.95
Yes 0.38 0.4 44.90 67.32 13.74 - - 41.98
LM
& Code
No - - 41.47 - - 35.98 33.00 36.82
[…]
Math
& Code
No - - - 65.05 10.37 9.15 9.80 23.59
[…]
@@ 3.2 Theoretical Analysis
According to the theoretical and empirical findings of Neyshabur, Sedghi, and Zhang (2020), fine-tuning models from the same pretrained initialization can guide them into a shared, flat, and connected low-loss basin. As a result, the linear interpolation path between such fine-tuned models remains within this low-loss basin and does not encounter significant performance barriers. In some cases, models along the interpolation path even outperform the endpoints.
@@
$f_{i}(\lambda_{m})=\lambda_{m}\,\delta_{t1,i}+(1-\lambda_{m})\,\delta_{t2,i},$ (11)
@@ 4.1 Experimental Setup
Task-Specific Fine-Tuned LLMs and Datasets.
Following the experimental setup given in Yu et al. (2024), we select three task-specific fine-tuned LLMs: WizardLM-13B (Xu et al. 2024), WizardMath-13B (Luo et al. 2023), and llama-2-13b-code-alpaca (Chaudhary 2023), all of which use Llama-2-13b (Touvron et al. 2023) as the pre-trained backbone. These models are respectively designed for instruction-following, mathematical reasoning, and code generation tasks. To evaluate the instruction-following task we use AlpacaEval (Li et al. 2023). For testing mathematical reasoning task, we employ GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021). For estimating the performance of code-generating task, we use HumanEval (Chen et al. 2021) and MBPP (Austin et al. 2021). More details of these LLMs and datasets can be found in Appendix A.
Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness.
To evaluate OOD robustness, we evaluate merged models including math & code, LM & math, and LM & code on instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al. 2024), respectively. More details on OOD benchmarks are given in Appendix B.
We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al. 2024) to evaluate the robustness of LLMs against adversarial prompts. Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al. 2018), BERTAttack (word-level) (Li et al. 2020), and StressTest (sentence-level) (Naik et al. 2018). The evaluation is conducted on two datasets supported by PromptBench: SST2 (sentiment analysis) (Socher et al. 2013) and CoLA (grammatical correctness) (Warstadt 2019). For more details on PromptBench and attack methods, please refer to Appendix C.
Evaluation Metrics.
@@
Unless otherwise specified, the remaining details of the model merging experiments follow Yu et al. (2024). For a detailed description of the hyperparameter settings used in the representative existing model merging methods, please refer to Appendix D. All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
Model Dataset
Use
Mixup
Use
Attack
Metric (%) PDR (%)
Math
Code
SST2 No No 57.68 38.97